OpenStudy (anonymous):

The isotope Cobalt-60 has a nuclear mass of 59.933820 u Calculate the Mass Defect of Cobalt-60 using the following information. Mass of Proton: 1.007825 u Mass of Neutron: 1.008665 u 1 u = 931.5 MeV

OpenStudy (anonymous):

What we want to do is figure out what the mass of the nucleus' particles are when seperated, and compare it to the mass of the actual nucleus. Cobalt has atomic number 27 which means it has 27 protons. The 60 means the amount of protons plus neutrons equals 60, so there are 33 neutrons. \[Mass of 27 protons=27 \times 1.007825u = 27.211275u\] \[Mass of 33 neutrons=33 \times 1.008665=33.285945u\] \[Total=60.49722\] The difference between this and the actual mass (59.933820u) is the mass defect, which is .5634u. Convert to MeV- \[.5634u \times 931.5MeV/u=524.8071 MeV\]

OpenStudy (anonymous):

Where did you get the 0.5634 u from?

OpenStudy (anonymous):

Also the 931.5MeV/u ?? Where did those numbers come from

OpenStudy (anonymous):

Oh nvm I see where you got the 931 one from but not the 0.5634

OpenStudy (anonymous):

.5634 is the actual mass of the nucleus (59.933820) subtracted from the total mass of its parts (60.49722)

OpenStudy (anonymous):

So you took 60.49722 - 59.933820

OpenStudy (anonymous):

Right. Thats the mass defect. Then I just converted it using the conversion factor given in the question